OpenStudy (anonymous):

Find if no solution, only one solution and if so, find the variables, or infinitely many solutions? -4x+6y-5z=-5 -2x+5y-3z=1 3x-2y+3z=3

OpenStudy (campbell_st):

the 1st task is to reduce the 3 equations to 2 equations with 2 unknowns Equation number -4x + 6y - 5z = -5 (1) -2x + 5y - 3z = 1 (2) 3x - 2y + 3z = 3 (3) I'd choose to eliminate x so its equation 1 + -2 * equation 2 -4x + 6y - 5z = -5 4x - 10y +6z = -2 ------------------- -4y + z = -7 (4) next 3* equation (2) + 2* equation (3) -6x + 15y - 9z = 3 6x - 4y + 6z = 6 -------------------- 11y -3z = 9 (5) now you have 2 equations -4y + z = -7 (4) 11y - 3z = 9 (5) eliminate z 3*equation (4) + equation (5) -12y + 3z = -21 11 y - 3z = 9 then -y = -12 or y = 12 substitute into equation (4) to find z -4(12) + z = -7 - 48 + z = -7 z = 41 now with y =12 and z = 41 substitute into equation 1 to find x -4x + 6(12) - 5(41) = -5 -4x +72 - 205 = -5 x = -32 so the solution is x = -32, y = 12 and z = 41